10. Calculate the number of molecules of sulphur (S8) present in 16 g of solid sulphur.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
1 mole of solid sulphur (S8) = 8 × 32 g = 256 g i.e., 256 g of solid sulphur contains = 6.022 × 1023 molecules Then, 16 g of solid sulphur contains = 6.022 × 1023/256 × 16 molecules = 3.76375 × 1022 molecules